# Spec-Gloss Surfels and Normal–Diffuse Priors for Relightable Glossy Objects

### v1

@@ Acknowledgments
This work is supported by the Flanders AI Research Program. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
@@ 6 Indirect Illumination
We adopt the indirect lighting formulation of Ref-Gaussian [33]. The specular IBL is split into direct and indirect components, modulated by visibility along the reflected direction $\mathbf{\omega}_{r}=2(\mathbf{\omega}_{o}\cdot\mathbf{n})\mathbf{n}-\mathbf{\omega}_{o}$ that is computed via bounding volume hierarchy (BVH) for accelerated ray tracing on a periodically extracted TSDF mesh. The visible part uses the standard prefiltered environment lookup, while the occluded part is modeled per Gaussian with low-order spherical harmonics $S_{i}(\cdot)$ evaluated at the reflected direction and alpha-blended in screen space:
$L_{\text{ind}}(\mathbf{p})=\sum_{i=1}^{N}l_{\text{ind}}\,\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}),$ (11)
where $l_{\text{ind}}=S_{i}(\mathbf{p},\mathbf{n})$. This formulation captures inter-reflections while retaining real-time performance.
@@ 7 Material Parameterizations in the Literature
which entangles the base color $b$ to the metallic $m$ through $F_{0}$. GaussianShader [9] simplifies the rendering equation to solve the entanglement issue by learning $F_{0}$ and using it for energy conservation on the diffuse appearance as well, departing, however, from principled models in literature or industry.
@@
where $\sigma$ is the sigmoid function. More recently, Ref-Gaussian [33] tried a similar approach that solved the entanglement issue by replacing base color $b$ in $F_{0}$ with a specular tint term $k_{s}$
$L_{o}^{\text{MSR}}=b+(F_{0}(m,k_{s})\beta_{1}+\beta_{2})L_{s},$ (14)
where, the Fresnel coefficients $\beta_{1}$, $\beta_{2}$ are given from a lookup table indexed with roughness $r$ and the $nv$ product of the surface normal $\mathbf{n}$ and viewing ray direction $\mathbf{v}$. Finally, $L_{d}$ is the irradiance from the environment, convolved with a diffuse kernel, and $L_{s}$ is the prefiltered specular environment. Nevertheless, it faces the same portability issues as GaussianShader, and while both methods perform well on the NVS task, performance deteriorates in relighting.
@@ 8 Additional Results
8.1 Relighting
Fig. 12 and Tab. 5 compare relighting results by our method and other Gaussian Splatting baselines. Our method generates more plausible relighting compared to Ref-Gaussian [33], 3DGS-DR [12], and GaussianShader [9]. Fig. 13 compares relighting results among the three material parameterizations (SG, MR, MRS) that were previously discussed. This figure validates our selection of the Spec-Gloss material parameterization, since it consistently outperforms the MR and MRS variants.
Figure 12: Qualitative comparison of relighting across three distinct environment maps (corridor, golf, neon) on the Glossy Synthetic dataset.
Figure 13: Qualitative comparison of relighting across three distinct environment maps (corridor, golf, neon) on the Glossy Synthetic dataset.
Table 5: Per-scene relighting and NVS comparison on the Glossy Synthetic [21] scenes. We highlight the first, second, and third best results.
@@ 8.1 Relighting
GShader [9] 20.34 15.92 15.83 14.04 18.47 19.25 17.31 28.07 31.81 27.18 30.09 24.48 23.58 27.55
3DGS-DR [12] 20.59 20.36 21.34 20.01 19.08 19.94 20.22 31.65 33.86 28.71 32.79 28.94 25.36 30.14
RefGauss [33] 19.87 21.58 20.45 20.20 21.35 21.45 20.82 32.86 33.01 30.04 33.07 29.84 26.68 30.92
@@
GShader [9] 0.900 0.868 0.828 0.783 0.875 0.883 0.856 0.919 0.961 0.914 0.938 0.898 0.901 0.921
3DGS-DR [12] 0.886 0.894 0.875 0.862 0.883 0.906 0.884 0.962 0.976 0.936 0.957 0.952 0.936 0.953
RefGaass [33] 0.888 0.919 0.877 0.870 0.912 0.894 0.893 0.969 0.973 0.952 0.963 0.962 0.947 0.961
@@
GShader [9] 0.114 0.104 0.098 0.156 0.138 0.115 0.121 0.098 0.056 0.064 0.088 0.122 0.091 0.086
3DGS-DR [12] 0.103 0.097 0.083 0.115 0.110 0.080 0.098 0.064 0.040 0.053 0.075 0.067 0.067 0.058
RefGauss [33] 0.104 0.070 0.079 0.110 0.087 0.087 0.090 0.040 0.040 0.043 0.064 0.058 0.058 0.051
@@ 8.2 Recovered Environment Map Results
Figures 14-16 present additional results on environment map recovery for the Shiny Synthetic [30], Glossy Synthetic [21], and Shiny Real [30] scenes.
Figure 14: Comparison of recovered environment maps.
Figure 15: Comparison of recovered environment maps.
Figure 16: Comparison of recovered environment maps.
@@ 8.3 Reconstruction Results
In figures 17 - 19, we present qualitative results for all scenes in the Shiny Synthetic [30], Glossy Synthetic [21], and Shiny Real datasets [30]. The corresponding quantitative results are presented in Tab. 6. We observe that our method is competitive and often outperforms Gaussian splatting and implicit baselines. The gap in performance in real scenes is mostly due to the limitation of 2DGS in representing dense, thin structures and sparsely seen background features.
Figure 17: Comparison of reconstruction quality on the Shiny Synthetic dataset [30].
Figure 18: Comparison of reconstruction quality on the Glossy Synthetic dataset [21].
Figure 19: Comparison of reconstruction quality on the Shiny Real dataset [30]. In the leftmost column, we show the GT RGB test images and the pseudo-GT surface normals generated by StableNormal [34] and used to supervise the predicted and depth-gradient-based surface normals.
Table 6: Quantitative comparison of novel view synthesis averaged over scenes within each dataset. Methods labeled with ✓in the Rel. column support relighting while those labeled with ✗do not. For RefGauss [33], although higher scores were reported on Shiny Real [30], we were unable to reproduce them. We therefore report results obtained using the authors’ released code. We highlight the first, second, and third best results.
Method Rel. Shiny Synthetic [30] Glossy Synthetic [21] Shiny Real [30]
@@ 8.4 Material Decomposition
In Fig. 20 and Fig. 21, we present the material decomposition of our method, the diffuse and specular components, the surface normals, as well as visibility, direct and indirect light terms for each scene in the Shiny Synthetic [30] and Glossy Synthetic datasets [21].
Figure 20: Material decomposition and outputs of our method on the Shiny Synthetic dataset [30]
Figure 21: Material decomposition and outputs of our method on the Glossy Synthetic dataset [21]
8.5 StableDelight-based Diffuse Priors
Fig. 22 presents sample predictions of StableDelight [25] for all examined scenes from the Shiny Synthetic [30], Glossy Synthetic [21] and Shiny Real [30] scenes. StableDelight manages to remove specular reflections in most cases, but fails in cases of mirror-like surfaces, confusing the reflections as part of the diffuse appearance of the scene.
Figure 22: Example predictions of StableDelight [25] on scenes from the Shiny Synthetic [30], Glossy Synthetic [21], and Shiny Real [30] datasets.

### v2

@@ Acknowledgments
This work was supported by the Flanders AI Research Program and the KU Leuven Methusalem project ”Lifelines”. The resources and services used in this work were provided by the VSC (Flemish Supercomputer Center), funded by the Research Foundation - Flanders (FWO) and the Flemish Government.
[…]
@@ 6 Indirect Illumination
We adopt the indirect lighting formulation of Ref-Gaussian [yao2025refgaussian]. The specular IBL is split into direct and indirect components, modulated by visibility $V\in\{0,1\}$ along the reflected direction $\boldsymbol{\omega}_{r}=2(\boldsymbol{\omega}_{o}\cdot\mathbf{n})\mathbf{n}-\boldsymbol{\omega}_{o}$. The visibility is computed via bounding volume hierarchy (BVH) for accelerated ray tracing on an extracted TSDF mesh, which is updated periodically (i.e. every 3k iterations) for efficiency. The visible part uses the standard prefiltered environment lookup, while the occluded part is modeled per Gaussian with low-order spherical harmonics $S_{i}(\cdot)$ evaluated at the reflection direction $\boldsymbol{\omega}_{r}$ and alpha-blended in screen space:
$L_{\text{ind}}(\boldsymbol{\omega}_{r})=\sum_{i=1}^{N}l_{\text{i}}(\boldsymbol{\omega}_{r})\,\alpha_{i}\ \prod_{j=1}^{i-1}(1-\alpha_{j}),$ (9)
where $l_{\text{i}}(\boldsymbol{\omega}_{r})=S_{i}(\boldsymbol{\omega}_{r})$. This formulation captures inter-reflections while retaining real-time performance.
@@ 7 Material Parameterizations in the Literature
which entangles the base color $b$ to the metallic $m$ through $F_{0}$. GaussianShader [jiang2024gshader] simplifies the rendering equation to solve the entanglement issue by learning $F_{0}$ and using it for energy conservation on the diffuse appearance as well, departing, however, from principled models in literature or industry. The simplified outgoing radiance equation is defined as:
@@
where $\sigma$ is the sigmoid function. More recently, Ref-Gaussian [yao2025refgaussian] tried a similar approach that solved the entanglement issue by replacing base color $b$ in $F_{0}$ with a specular tint term $k_{s}$
$L_{o}^{\text{MRS}}=b+(F_{0}(m,k_{s})\beta_{1}+\beta_{2})L_{s},$ (12)
where the Fresnel coefficients $\beta_{1}$, $\beta_{2}$ are given from a lookup table indexed with roughness $r$ and the nvnv product of the surface normal $\mathbf{n}$ and viewing ray direction $\mathbf{v}$. Finally, $L_{d}$ is the irradiance from the environment, convolved with a diffuse kernel, and $L_{s}$ is the prefiltered specular environment. Nevertheless, it faces the same portability issues as GaussianShader, and while both methods perform well on the NVS task, performance deteriorates in relighting.
@@ 8 Additional Results
8.1 Scene Editing
Tabs. 5 and 12 compare our relighting results with Gaussian Splatting baselines, showing that our method produces more plausible relighting than Ref-Gaussian [yao2025refgaussian], 3DGS-DR [ye2024gsdr], and GaussianShader [jiang2024gshader]. Fig. 13 compares the three material parameterizations (SG, MR, MRS) and validates our choice, as SG consistently outperforms the MR and MRS variants. Fig. 14 further illustrates the material editing capabilities of our method.
Table 5: Per-scene relighting and NVS comparison on the Glossy Synthetic [liu2023nero] scenes. We highlight the first, second, and third best results.
@@ 8.1 Scene Editing
GShader [jiang2024gshader] 20.34 15.92 15.83 14.04 18.47 19.25 17.31 28.07 31.81 27.18 30.09 24.48 23.58 27.55
3DGS-DR [ye2024gsdr] 20.59 20.36 21.34 20.01 19.08 19.94 20.22 31.65 33.86 28.71 32.79 28.94 25.36 30.14
RefGauss [yao2025refgaussian] 19.87 21.58 20.45 20.20 21.35 21.45 20.82 32.86 33.01 30.04 33.07 29.84 26.68 30.92
@@
GShader [jiang2024gshader] 0.900 0.868 0.828 0.783 0.875 0.883 0.856 0.919 0.961 0.914 0.938 0.898 0.901 0.921
3DGS-DR [ye2024gsdr] 0.886 0.894 0.875 0.862 0.883 0.906 0.884 0.962 0.976 0.936 0.957 0.952 0.936 0.953
RefGauss [yao2025refgaussian] 0.888 0.919 0.877 0.870 0.912 0.894 0.893 0.969 0.973 0.952 0.963 0.962 0.947 0.961
@@
GShader [jiang2024gshader] 0.114 0.104 0.098 0.156 0.138 0.115 0.121 0.098 0.056 0.064 0.088 0.122 0.091 0.086
3DGS-DR [ye2024gsdr] 0.103 0.097 0.083 0.115 0.110 0.080 0.098 0.064 0.040 0.053 0.075 0.067 0.067 0.058
RefGauss [yao2025refgaussian] 0.104 0.070 0.079 0.110 0.087 0.087 0.090 0.040 0.040 0.043 0.064 0.058 0.058 0.051
@@
Figure 12: Qualitative relighting comparison under three environment maps (corridor, golf, neon) on the Glossy Synthetic dataset. Our method typically yields more faithful relighting than competing Gaussian Splatting baselines.
Figure 13: Qualitative relighting comparison under three environment maps (corridor, golf, neon) on the Glossy Synthetic dataset. We compare our method against three alternative material parameterizations, and the results support our choice of the specular–glossiness (SG) parameterization.
Figure 14: Demonstration of the material-editing capabilities of our method. After training, we modify the reconstructed materials and re-render three scenes from the Shiny Synthetic dataset [verbin2022refnerf] with increased or decreased albedo, roughness, or $F_{0}$.
@@ 8.2 Recovered Environment Map Results
Figures 15-17 present additional results on environment map recovery for the Shiny Synthetic [verbin2022refnerf], Glossy Synthetic [liu2023nero], and Shiny Real [verbin2022refnerf] scenes.
Figure 15: Comparison of recovered environment maps on the scenes of the Shiny Synthetic dataset [verbin2022refnerf].
Figure 16: Comparison of recovered environment maps from the scenes of the Glossy Synthetic dataset [liu2023nero].
Figure 17: Comparison of recovered environment maps on the scenes of the Shiny Real dataset [verbin2022refnerf].
@@ 8.3 Reconstruction Results
In Figures 18 - 20, we present qualitative results for all examined scenes. The corresponding quantitative results are presented in Tab. 6. We observe that our method is competitive and often outperforms Gaussian splatting and implicit baselines. The gap in performance in real scenes is mostly due to the limitation of 2DGS in representing dense, thin structures and sparsely seen background features.
Figure 18: Comparison of rendering quality on the Shiny Synthetic dataset [verbin2022refnerf].
Figure 19: Comparison of rendering quality on the Glossy Synthetic dataset [liu2023nero].
Figure 20: Comparison of rendering quality on the Shiny Real dataset [verbin2022refnerf]. In the leftmost column, we show the GT RGB test images and the pseudo-GT surface normals, generated by StableNormal [ye2024stablenormal], and used to supervise the predicted normals.
Table 6: Quantitative comparison of novel view synthesis averaged over scenes within each dataset. Methods labeled with ✓in the Rel. column support relighting while those labeled with ✗do not. For RefGauss [yao2025refgaussian], although higher scores were reported on Shiny Real [verbin2022refnerf], we were unable to reproduce them. We therefore report results obtained using the authors’ released code. We highlight the first, second, and third best results.
Method Rel. Shiny Synthetic [verbin2022refnerf] Glossy Synthetic [liu2023nero] Shiny Real [verbin2022refnerf]
@@ 8.4 Material Decomposition
In Fig. 21 and Fig. 22, we present the material decomposition of our method, the diffuse and specular components, the surface normals, as well as visibility, direct and indirect light terms for all examined synthetic scenes.
Figure 21: Material decomposition and outputs of our method on the Shiny Synthetic dataset [verbin2022refnerf]
Figure 22: Material decomposition and outputs of our method on the Glossy Synthetic dataset [liu2023nero]
8.5 Diffuse Color and Surface Normal Priors
Fig. 23 and Fig. 24 present sample predictions of StableDelight [stabledelight2025] and StableNormal [ye2024stablenormal] for all examined scenes from the Shiny Synthetic [verbin2022refnerf], Glossy Synthetic [liu2023nero] and Shiny Real [verbin2022refnerf] scenes. StableDelight manages to remove specular reflections in most cases, but fails in cases of mirror-like surfaces (e.g. toaster and glossy teapot), confusing the reflections as part of the diffuse appearance of the scene. StableNormal, on the other hand, fails to recognize the surface of the liquid in the coffee scene and oversmooths complex geometries (e.g. cat, luyu) and occasionally confuses interreflections as separate geometries (e.g. tbell).
Figure 23: Example predictions of StableDelight [stabledelight2025] on scenes from the Shiny Synthetic [verbin2022refnerf], Glossy Synthetic [liu2023nero], and Shiny Real [verbin2022refnerf] datasets.
Figure 24: Example predictions of StableNormal [ye2024stablenormal] on scenes from the Shiny Synthetic [verbin2022refnerf], Glossy Synthetic [liu2023nero], and Shiny Real [verbin2022refnerf] datasets.